OpenStudy (anonymous):

How would I solve a problem like this? Sarah has a square garden with an area of ______ square feet. She wants to put a chain link fence along one complete side of her garden. Fencing is sold in whole numbers of feet. What is the LEAST amount of fencing she will have to buy?

OpenStudy (anonymous):

@campbell_st @Embryo

OpenStudy (anonymous):

Oh my, i missed the part where it said that it was a square garden, that makes it a lot simpler, so because it's a square garden, if the\[Area=x\]do you know what the length of one side would be?

OpenStudy (anonymous):

1/4 of the area? @Embryo

OpenStudy (anonymous):

@jim_thompson5910 @ranga

OpenStudy (anonymous):

Not quite, that would be the length of one side if you had the perimeter, remember\[Perimeter=4s\]\[Area=s^2\]

OpenStudy (anonymous):

What does S stand for?

OpenStudy (anonymous):

Oh sorry, so "S" stands for the length of one side

OpenStudy (anonymous):

OK, so the side times itself?

OpenStudy (anonymous):

That's how you Calculate area yes, so back to the original thing, you have a square with area=X square feet, if that's true, what was the length of one of the sides that made that true?

OpenStudy (anonymous):

So if the problem was actually.. Myra has a square garden with an area of 52 square feet. She wants to put a chain link fence along one complete side of her garden. Fencing is sold in whole numbers of feet. What is the LEAST amount of fencing she will have to buy? A. 26 feet B. 13 feet C. 8 feet D. 7 feet

OpenStudy (anonymous):

You'd have to work backwards from the area you were given so, since you have\[Area=52=s^2\]this would imply that one of the sides was\[s^2=52 \rightarrow s=\sqrt{52}\]using the square root estimations you've been doing today, you need to round the square root of 52 upwards, since she wants to cover the ENTIRE fence

OpenStudy (anonymous):

7x7=49 So.... D?

OpenStudy (anonymous):

Not quite, remember, she wants to cover the ENTIRE side, so in order to do that, she has to round upwards, since 49<52, it wouldn't cover the entire side

OpenStudy (anonymous):

Oh! So it has to be at least 52?

OpenStudy (anonymous):

Yes, at least 52, but since\[\sqrt{52}\approx 7.2\]and she can only by the fence pieces in whole numbers, she has to buy 8 feet worth, see why? if she buys 7 feet, she won't cover the entire side

OpenStudy (anonymous):

Yes!!! I get it!!! :p Thank you so much!!! :)